can you help me set up this problem and solve
the current I in electrical conductor varies inversely as the resistance R of the conductor. suppose I is 0.2 amperes when the resistance is 200ohms. find the current when the resistance is 250ohms.

Answer by ankor@dixie-net.com(22740)

the current I in electrical conductor varies inversely as the resistance R of the conductor. suppose I is 0.2 amperes when the resistance is 200ohms. find the current when the resistance is 250ohms.
:
Use Ohm's law to find the voltage: E = I*R
E = .2*200
E = 40 v
:
I = E/R
When 40 volts is applied to a resistance of 250 ohms
I = 40/250
I = .16 amps